# Deepfake Detection that Generalizes Across Benchmarks

### v1

@@ 4.3 Ablation studies
We conducted ablation experiments to understand the contribution of each component, see  Tab. 3. There are five setups. Setup 5 is the proposed LNCLIP-DF:
@@
    Baseline: the setup is similar to [35]. We train only a linear classifier on top of the features given by the CLIP ViT’s classification token. The ViT remains frozen.
@@
    Baseline + LN + L2: adds L2 normalization of the classification token to setup 2.
  4. 4.
    Baseline + LN + L2 + UnAl: adds uniformity and alignment losses to setup 3.
  5. 5.
    Baseline + LN + L2 + UnAl + slerp: adds slerp augmentations to setup 4.
Table 5: Ablation study. Impact of model components on test video-level AUROC in cross-dataset settings. Each setup was trained on the FF++ training set. Setup #5 is the proposed LNCLIP-DF.
# Setup UADFV DFD DFDC FSh CDFv2 FFIW KoDF FAVC DFDM PGF IDF DSv1.1 DSv2 Mean
1 Baseline 95.75 86.7 74.0 77.88 78.8 84.6 82.2 80.8 80.7 61.1 68.17 61.6 56.6 76.1
2 #1 + LN 99.25 96.6 84.0 86.66 93.1 91.6 87.4 94.8 99.1 87.8 93.32 89.1 77.7 90.8
3 #2 + L2 99.83 97.3 86.6 88.87 94.5 91.6 84.5 96.0 99.4 89.6 95.17 89.8 79.8 91.8
4 #3 + UnAl 99.63 97.8 86.5 87.32 94.3 90.7 87.0 96.1 99.6 90.7 97.97 90.6 78.6 92.1
5 #4 + slerp 99.83 97.3 87.0 88.22 96.5 93.5 85.9 96.3 98.8 92.9 97.30 92.6 80.0 92.8
Layer Normalization tuning.
As shown in Tab. 3, LN-tuning consistently achieves substantial gains in AUROC across all benchmarks.
@@
L2 normalization.
L2 normalization of the classification token results in a more stable increase in the validation AUROC and a slower overfitting during training. The single change improves generalization, as shown by the performance difference between setups #2 and #3 in Tab. 5.
Uniformity and Alignment.
We hypothesized that representations that effectively exploit the latent space could help us improve generalization. We used uniformity and alignment loss [48], leading to slower overfitting and better validation AUROC. We suppose that adding these losses helps to utilize the hyperspherical space more effectively, resulting in improved generalization.
Slerp.
Adding slerp augmentation, described in Sec. 3.1, improves average test AUROC. We hypothesize that slerp promotes generalization by covering the feature space more densely, filling the gaps between features of the same class. We believe that by enriching the feature space with the interconnection of latent features, slerp prepares the model for recognizing signatures of fake samples from unseen generation techniques whose latent representations can resemble combinations of previously seen forgery signatures.
@@ 4.4 Importance of training on paired dataset
In particular, we generate 10 paired and 10 unpaired datasets from the FF++ training set. In each of the 10 trials, the training set is randomly split into two halves. For each paired dataset, we include all real videos from the first half along with deepfakes generated from those videos. For each unpaired dataset, we use the same real from the first half and deepfake videos from the second half, thus ensuring that none of them is generated from the corresponding real video.
We plot AUROC for training and validation progress in Fig. 3. Models trained on the unpaired datasets consistently begin to overfit more quickly and achieve a lower validation AUROC, as indicated by the sharper increase in training and the decline in the validation AUROC. In contrast, models trained using paired datasets achieve a higher validation AUROC and have a lower tendency to overfit throughout the training process.
@@
Figure 3: Video-level AUROC for (a) Training and (b) Validation, averaged over 10 randomly sampled paired and unpaired datasets.
@@ 5 Limitations and Future Work
Although the proposed method, LNCLIP-DF, demonstrates a strong generalization, it is important to acknowledge its limitations, which also present avenues for future research.
Robustness to perturbations and adversarial attacks.
One limitation is that the model’s performance against targeted adversarial attacks [45] specifically designed to evade detection has not been evaluated. Furthermore, while standard image augmentations were used in training, a more systematic analysis of the model’s robustness to a wider range of unseen data perturbations and quality degradations is necessary.
Limited temporal modeling.
The proposed method operates at the frame level and aggregates predictions via simple averaging. Although this allows for architectural simplicity, it ignores temporal dynamics that could offer additional forgery cues, such as inconsistencies in lip synchronization or frame-level motion artifacts.
Dependence on facial region.
Our preprocessing pipeline assumes access to a detectable and alignable face. Videos with occlusions, extreme poses, or low resolution may lead to failures in face detection or misaligned crops, which can degrade performance.
Limited demographic diversity in training data.
our work provides several key insights for the community Training data lack sufficient diversity in terms of identity, ethnicity, age, and gender. This can result in biased detection performance, where the model may generalize poorly to underrepresented demographic groups. Ensuring demographic balance and evaluating fairness across subgroups are important next steps toward responsible deployment.
Incremental learning.
The current framework does not address the challenge of incremental learning. As new forgery methods emerge, the model would require complete retraining. Developing a framework that can adapt to new manipulation techniques without catastrophically forgetting previously learned ones is a critical direction for real-world applicability.
@@ 6 Conclusions
In this work, we addressed the persistent challenge of generalization in deepfake detection. Although recent approaches have explored increasingly complex architectural and parameter-space modifications to adapt large foundation models, we show that generalization is achieved by tuning existing parameters of the CLIP image encoder. We demonstrated that CLIP can be transformed into a state-of-the-art deepfake detection model using four additional components: LN-tuning, L2 normalization, slerp augmentations, and a linear classifier. State-of-the-art results can be achieved within just a few hours of training on a single A100 GPU, making the proposed approach computationally efficient and reproducible.

### v2

@@ 4.3 Ablation studies
We conducted ablation experiments to understand the contribution of each component; see Tab. 6. There are three setups tested for every backbone. Setup 3 is the proposed GenD:
@@
    Baseline: the setup is similar to [40]. We freeze the backbone and train a linear classifier with cross-entropy loss on top of the feature space of the classification token.
@@
    Baseline + LN + UA: L2-normalizes classification token and adds uniformity and alignment losses to setup 2.
Table 6: Ablation study. Impact of model components on test video-level AUROC in cross-dataset settings. Each setup was trained on the FF++ training set. Reported values are averages over 5 training seeds. Setups #3, #6, #9 are the proposed GenD.
# Setup UADFV DFD DFDC FSh CDFv2 FFIW KoDF FAVC DFDM PGF IDF DSv1.1 DSv2 CDFv3 Mean
1 CLIP + LP 94.6 89.2 75.3 77.6 74.6 80.7 81.8 83.0 77.8 62.7 68.7 64.5 57.8 75.6 76.0
2 #1 + LN 99.0 96.1 85.4 86.9 93.8 91.5 84.4 95.5 99.3 87.5 95.0 89.1 74.0 86.8 90.3
3 #2 + UA 99.2 96.4 86.4 86.6 94.6 91.5 84.9 96.0 99.6 89.6 97.8 90.1 77.7 85.9 91.2
4 $\text{PE}_{\text{core}}\text{L}$ + LP 76.1 76.4 67.2 80.3 72.2 75.7 65.6 71.3 73.3 52.8 74.6 65.6 59.5 79.5 70.7
5 #4 + LN 98.4 96.4 81.9 88.1 94.2 93.0 84.5 96.5 97.9 91.2 97.7 88.0 77.0 89.9 91.0
6 #5 + UA 97.7 96.8 82.2 87.6 95.0 93.7 85.1 97.3 98.3 92.3 97.9 87.8 78.6 89.5 91.4
7 DINOv3L + LP 84.2 79.8 69.0 71.9 68.2 74.7 75.4 81.1 85.9 65.0 82.0 72.2 70.2 68.3 74.9
8 #7 + LN 97.3 94.8 84.4 88.7 91.7 90.4 90.6 98.4 98.8 93.6 97.5 89.1 82.4 80.1 91.3
9 #8 + UA 98.6 96.2 85.6 88.8 92.5 92.9 89.7 98.4 99.8 92.4 98.2 86.9 79.4 83.5 91.6
Layer Normalization tuning. As shown in Tab. 6, LN-tuning consistently achieves substantial gains in AUROC across all benchmarks.
@@
Uniformity and Alignment. We hypothesized that representations that effectively exploit the latent space could help us improve generalization. We used uniformity and alignment loss [52], leading to slower overfitting and better validation AUROC. We suppose that adding these losses helps to utilize the hyperspherical space more effectively, resulting in improved generalization.
@@ 4.4 Importance of training on paired dataset
In particular, we generate 20 different paired and 20 unpaired dataset splits from the FF++ training set. Every split shares the real part but has a different fake part. In the paired dataset, every fake video shares the background with a video from the real part. In an unpaired dataset, every fake video does not have a corresponding real video with the same background, see Fig. 2. We ensure that the number of fake and real videos between the paired and unpaired datasets is the same.
We plot AUROC for training and validation progress of $\text{PE}_{\text{core}}\text{L}$ image encoder in Fig. 3. Models trained on the unpaired datasets consistently overfit more quickly and achieve a lower validation AUROC, as indicated by the sharper increase in training and the decline in the validation AUROC. In contrast, models trained using paired datasets achieve a higher validation AUROC and have a lower tendency to overfit throughout the training process. We observe the same behavior on other pre-trained backbones and training sets; see supplementary material for more experiments.
@@
Figure 3: Video-level AUROC for (a) Training and (b) Validation, averaged over 20 randomly sampled paired and 20 unpaired datasets from FF++. The image encoder is $\text{PE}_{\text{core}}\text{L}$.
@@ 4.5 Evolution of detection difficulty over the years
4.6 Robustness to image degradations
We evaluated the robustness of the GenD to commonly used image degradations, which can be used to remove forgery patterns and decrease the performance of deepfake detection. For statistical significance, we average the results of GenD trained using 5 different training seeds. The results are shown in Fig. 5. Visual examples of images processed over various levels of degradations can be found in the supplementary.
We selected three methods, such as Gaussian blur, JPEG compression, and image resizing. For Gaussian blur, we varied the sigma $\sigma$ and selected the proportional kernel size $k=2\cdot\left\lceil 3\sigma\right\rceil+1$. For resizing, we took the facial image and resized it to $224^{2}$, $114^{2}$, or $64^{2}$ pixels by using 6 different interpolation techniques: nearest, lanczos, bilinear, bicubic, box, hamming. After that, each downsized image was upsampled to the model’s input resolution by every method’s preprocessing technique. Results are averaged over the interpolation methods.
@@ 4.6 Robustness to image degradations
Figure 5: Robustness to image degradations for GenD ($\text{PE}_{\text{core}}\text{L}$), ForAda [13], and Effort [58]. Video-level AUROC (%) is calculated across all 14 test datasets. Error bars for GenD are computed from models trained with 5 different training seeds. In the resize, we also average every method across 6 interpolation strategies: nearest, lanczos, bilinear, bicubic, box, hamming.
@@ 5 Limitations and Future Work
Although the proposed method, GenD, demonstrates a strong generalization, it is important to acknowledge its limitations, which also present avenues for future research.
Robustness to perturbations and adversarial attacks. One limitation is that the model’s performance against targeted adversarial attacks [49] specifically designed to evade detection has not been evaluated.
Limited temporal modeling. We tried to model the temporal signal using a simple single-layer self-attention block between classification tokens from different frames in a video. Nevertheless, our experiments did not show any noticeable improvements compared to the proposed method, which operates at the frame level and aggregates predictions via simple averaging. Although this allows for architectural simplicity, it ignores temporal dynamics that could offer additional forgery cues, such as inconsistencies in lip synchronization or frame-level motion artifacts.
Dependence on facial region. Our preprocessing pipeline assumes access to a detectable and alignable face. Videos with occlusions, extreme poses, or low resolution may lead to failures in face detection or misaligned crops, which can degrade performance.
Limited demographic diversity in training data. Training data lacks sufficient diversity in terms of identity, ethnicity, age, and gender. This can result in biased detection performance, where the model may generalize poorly to underrepresented demographic groups. Ensuring demographic balance and evaluating fairness across subgroups are important next steps toward responsible deployment. See the supplementary for typical failure cases.
Incremental learning. The current framework does not address the challenge of incremental learning. As new forgery methods emerge, the model would require complete retraining. Developing a framework that can adapt to new manipulation techniques without catastrophically forgetting previously learned ones is a critical direction for real-world applicability.
@@ 6 Conclusions
We addressed the persistent challenge of generalization in deepfake detection. Although recent approaches explored increasingly complex architectural and parameter-space modifications to adapt large foundation models, we show that generalization is achieved by tuning existing parameters of the image encoder (CLIP, PE, DINO). We demonstrated that such encoder is transformed into a state-of-the-art deepfake detection model using three additional components: LN-tuning, L2 normalization, and a linear classifier. State-of-the-art results are achieved within a few hours of training on a single A100 GPU, making the proposed approach computationally efficient and reproducible.